# Describing the speed of sound peak of isospin-asymmetric cold strongly interacting matter using effective models

### v1

@@ V Summary and discussion
Our results show that an interesting way of reconciling results for the non-monotonic behavior of $c_{s}^{2}$ in isospin imbalanced strongly interacting matter, obtained with effective models and LQCD, may be closely connected with the introduction of medium-dependent couplings. In this work, we have determined the dependence of the couplings with $\mu_{I}$ using as input the LQCD data for the isospin density $n_{I}$ [9] in a thermodynamic consistent way. The numerical results for the thermodynamic functions are consequently also in agreement with LQCD results.
We find that the $\mu_{I}$ dependence of the couplings in the NJL and LSMq models turn out to have different behaviors. However, notice also that as opposed to the former, in the latter case we have only considered a medium-induced modification of $g$, ignoring possible medium-induced modifications of $\lambda$. The rationale behind this approximation rests on the assumption that since quantum effects are introduced in terms of a fermion loop [82], it is only the fermion coupling which should be considered to receive medium modifications. To try reconciling these behaviors, improvements to the approximations considered in LSMq calculations are currently being performed and will be reported elsewhere.

### v2

@@ 5 Summary and discussion
Our results show that an interesting way of reconciling results for the non-monotonic behavior of $c_{s}^{2}$ in isospin imbalanced strongly interacting matter, obtained with effective models and LQCD, may be closely connected with the introduction of medium-dependent couplings. In this work, we have determined the dependence of the couplings with $\mu_{I}$ using as input the three-flavor LQCD data for the isospin density $n_{I}$ [9], and the two-flavor version of the NJL and LSMq. We verified that the introduction of a medium dependence through the coupling requires a careful analysis of the thermodynamics of the system, and the addition of a compensating function can recover the agreement between statistical mechanics definitions and thermodynamic relations [79]. As the isospin density was fitted to the lattice, the numerical results for the thermodynamic functions are consequently also in agreement with LQCD results. Again, our idea here is not to claim a perfect description, but to point out that medium contributions to the coupling, which are normally not accounted for in effective models, may be closely connected to the characterization of a non-monotonic behavior in the speed of sound. We have also argued that the use of two-flavor models to compare to a three-flavor LQCD result is justified since the effect of the strange quark on $\langle n_{I}\rangle$ is, in our effective model description, only captured at higher orders which are suppressed with respect to the leading order calculation shown in this work. Nevertheless, it is important to mention that we have results from a preliminary analysis, using the the three-flavor NJL model of Ref. [89], to explore whether the description of this effect is different from what we obtain with the two-flavor models. What we find is that the results for the $G$-coupling are very similar in the two cases, showing the same overall shape as a function of $\mu_{I}$, decreasing towards a minimum and increasing afterwards. The slightly different values can then be attributed to the effect of strangeness, verifying that this has quantitatively a small effect. The similarities reflect the fact that, although the increase of degrees of freedom in the effective model may correspond to a fairer description of LQCD data, the two-flavor models already capture the physics encoded in the behavior of the coupling that is in turn responsible for the peak in the speed of sound. This result is part of an ongoing effort and will be complemented with a corresponding three-flavor analysis in the LSMq, which is more involved and will be reported elsewhere.
We find that the $\mu_{I}$ dependence of the couplings in the NJL and LSMq models turn out to have different behaviors. However, notice also that as opposed to the former, in the latter case we have only considered a medium-induced modification of $g$, ignoring possible medium-induced modifications of $\lambda$. The rationale behind this approximation rests on the assumption that since quantum effects are introduced in terms of a fermion loop [90], it is only the fermion coupling which should be considered to receive medium modifications. To try reconciling these behaviors, improvements to the approximations considered in LSMq calculations are currently being performed and will also be reported elsewhere.